# HyperSIGMA: Hyperspectral Intelligence Comprehension Foundation Model

### v1

@@ 5.6.1 Experiment Settings
Implementation Details The architecture for performing HSI super-resolution is similar to HSI denoising, as shown in Fig. 20. The main differences are: (1) the patch size of the spatial subnetwork is set to 1, and (2) after the reconstruction decoder, we use PixelShuffle [233] to upsample the reconstruction result (e.g., a scale factor of 8$\times$ corresponds to upsampling 3 times). This upsampling operation ensures the network outputs the expected size.
In our experiments, we first acquire training patches of size $Ms\times Ms$. These patches are uniformly downsampled to $s\times s$ using bicubic interpolation. We then upsample the downsampled patches back to the original size ($Ms\times Ms$) using various super-resolution methods. Specifically, we compare our SpatSIGMA and HyperSIGMA with several advanced methods, including Bicubic, SwinIR [228], GDRRN [227], SSPSR [229], RFSR [230], GELIN [231], and MSDFormer [232]. We examine two different $M$ settings, 4 and 8, while keeping $s$ fixed at 32.
We evaluate the models using six popular metrics to comprehensively assess performance in both spatial and spectral dimensions: Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM), Spectral Angle Mapper (SAM), Cross Correlation (CC), Root-Mean-Squared Error (RMSE), and Erreur Relative Globale Adimensionnelle de Synthèse (ERGAS). While PSNR, SSIM, and RMSE are standard metrics for natural image restoration, CC, SAM, and ERGAS are common in hyperspectral image fusion tasks. Superior super-resolution performance is indicated by higher PSNR, SSIM, and CC values, and lower SAM, RMSE, and ERGAS values.
5.6.2 Results and Analyses
The quantitative results of HSI super-resolution are shown in Table XIV, with visualization comparisons, including zoomed-in details, presented in Fig. 22. Table XIV reveals that our proposed HyperSIGMA outperforms other methods at both scale factors. Notably, HyperSIGMA surpasses the recent advanced approach, MSDFormer [232], across all metrics. Fig. 22 illustrates the effectiveness of our methods at an 8$\times$ scale factor, showing clearer visual results compared to other competitors, particularly in the highlighted zoomed-in areas. These findings further demonstrate the capability of our models in HSI low-level tasks.
5.7 Further Investigations and Analyses
5.7.1 Model Scalability
TABLE XV: Fine-tuning accuracies of SpatSIGMA and HyperSIGMA with different ViT backbones on Indian Pines and Xiongan datasets. Base: ViT-Base. Large: ViT-Large. Huge: ViT-Huge.
Method SpatViT SpecViT OA AA Kappa
Segmentation on Indian Pines
SpatSIGMA Base - 85.08 78.30 83.04
SpatSIGMA Large - 86.55 81.70 84.63
SpatSIGMA Huge - 86.74 75.02 84.89
HyperSIGMA Base Base 85.54 76.68 83.58
HyperSIGMA Base Large 86.81 79.49 85.00
HyperSIGMA Base Huge 87.38 83.24 85.63
HyperSIGMA Huge Huge 87.70 79.19 86.01
Classification on Xiongan
1DCNN - - 55.89 71.19 52.26
2DCNN - - 27.32 38.79 23.61
HybridSN - - 89.83 74.72 88.39
A2S2KResNet - - 90.32 76.85 88.93
SpatSIGMA Base - 93.68 79.80 92.75
SpatSIGMA Large - 94.59 82.95 93.78
SpatSIGMA Huge - 94.73 82.43 93.94
HyperSIGMA Base Base 95.06 82.43 94.32
HyperSIGMA Base Large 95.30 82.77 94.58
HyperSIGMA Base Huge 95.65 83.85 94.98
HyperSIGMA Huge Huge 95.78 84.19 95.13
Using Different ViT Backbones: In previous sections, we have shown that HyperSIGMA consistently outperforms state-of-the-art methods, even when using the ViT-Base backbone. To evaluate the scalability of HyperSIGMA, we fine-tuned SpatSIGMA and HyperSIGMA using different ViT backbones on the Indian Pines dataset and the more challenging Xiongan dataset [234], which is 1,580$\times$3,750 pixels with 256 channels and 19 categories, mainly comprising cash crops. We also implemented some classical methods for comparison, including 1DCNN and 2DCNN [30], HybridSN [235], and A2S2KResNet [236]. For Xiongan, models were trained for 300 epochs with a batch size of 512. The training set included 100 randomly selected samples per class, with the remainder used for evaluation. Experimental settings were consistent with those used in Sec. 5.1. Table XV presents the results. It is clear that using larger ViT backbones leads to better performance for both SpatSIGMA and HyperSIGMA, underscoring their good scalability. Notably, we observed that ViT-Huge in SpatSIGMA did not perform as well as lighter versions used in HyperSIGMA, highlighting the importance of utilizing spectral information. The largest HyperSIGMA, with over 1 billion parameters, achieved 87.70% OA and 95.78% OA on Indian Pines and Xiongan, respectively, surpassing other methods significantly.
@@ 5.7.1 Model Scalability
TABLE XVI: Classification accuracies of models using different pre-training weights on the HanChuan dataset.

### v2

@@ I.1 Implementation Details
Network Structure The architecture for performing HSI super-resolution is similar to HSI denoising, as shown in Fig. 35. The main differences are: (1) the patch size of the spatial subnetwork is set to 1, and (2) after the reconstruction decoder, we use PixelShuffle [219] to upsample the reconstruction result (e.g., a scale factor of 8$\times$ corresponds to upsampling 3 times). This upsampling operation ensures the network outputs the expected size.
Experimental Details All patches are uniformly downsampled to 32 $\times$ 32. Therefore, if for 8$\times$ super-resolution, we extract 256 $\times$ 256 patches.
I.2 More Results
Figure 37: Visualization of detected oil leakage in various Gulf of Mexico regions by our models. (a) GM07. (b) GM08. (c) GM09. (d) GM17.
Qualitative Results Fig. 36 illustrates the effectiveness of our methods at an 8$\times$ scale factor, showing clearer visual results compared to other competitors, particularly in the highlighted zoomed-in areas.
Other Upsampling Ratios We additionally conduct an experiment with a 4$\times$ super-resolution, as shown in Table XXIV. It can be seen that our models demonstrate clear performance advantages.
Appendix J Model Scalability Experiments
J.1 Implementation Details
Using Different ViT Backbones The experimental settings on the Indian Pines dataset are consistent with those used in Sec. D.1. For the Xiongan dataset, models were trained for 300 epochs with a batch size of 512. The training set included 100 randomly selected samples per class, with the remainder used for evaluation.
J.2 More results
@@ I.2 More Results
TABLE XXV: Classification accuracies of models using different pre-training weights on the HanChuan dataset.